GR014

PI: Rottgering
Title: Sample of dist gals

Session: Feb 1997 (session 1)
Frequency: 18 cm
Mode: B
Obs time (UT): EU 11:00(28/02)-21:00(28/02)
Stations: CbEbJbMcNtOnTrSmShUrWb+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Tugnoli Marco at Fri Feb 28 16:15:29 EST 1997) Success

Onsala (Antonis Polatidis at Sat Mar 1 06:39:47 EST 1997) Success
No known problems. Low winds at last !!!!

Effelsberg (dgraham at Mon Mar 3 03:48:33 EST 1997) Success
-----------------------------OK------------------------------------, occasional short elevation drive problems, should not affect data

Westerbork (Tony Foley at Tue Mar 4 06:43:09 EST 1997) Success
Looks OK. Note that a different synthesizer was used to em24

Shanghai (Huang Xinyong at Wed Mar 5 05:15:41 EST 1997) Success
Clock offset = 0.54-0.56 msec. No problems, weather overcast.

Jodrell (Alastair Gunn at Tue Mar 18 12:38:33 MET 1997) Minor failures
Recording wrong polarisations until about 14:22. Missed pass 11, on source late at 14:44. Very bad interference.

Calibration (Fredrik T Rantakyro at Thu May 8 11:37:49 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name gr14.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Scott Aaron at Thu Apr 16 12:34:14 METDST 1998) Minor failures
Correlation completed. No tapes were recieved from Cambridge. Onsala and Urumqi observed the 18 cm session with swapped polarizations. Standard tracking problems with Torun.